# Maximum Likelihood Estimation Based Complex-Valued Robust Chinese Remainder Theorem and Its Fast Algorithm

### v1

@@ II-B Problem Description and SR-ADCs
Now we consider the problem of recovering complex-valued bandlimited signals using the C-CRT. For a single SR-ADC, one can compute the modulus of real numbers. Two combined SR-ADCs can obtain a modulus sampler with Gaussian integer $\mathsf{M}_{i}$, as illustrated in Fig. 3. For convenience, we denote $\mathsf{M}_{i}=\rho_{i}e^{\mathrm{i}\theta_{i}}$, where $\theta_{i}$ represents the angle of $\mathsf{M}_{i}$, and $\rho_{i}$ represents the dynamic range of the SR-ADC. Let $T$ be the sampling interval length of each SR-ADC and $\mathsf{f}_{k}=\mathsf{f}(kT)$. Then, the output $\mathsf{y}_{ki}$ can be expressed as
@@
By applying a phase shift, one can obtain $\mathsf{y}_{ki}$ through separately applying the modulo operation to its real and imaginary parts. To be specific, if we rewrite $\mathsf{y}_{ki}$ as $\mathsf{y}_{ki}=\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}+\mathrm{i}\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}$, then $\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}$ and $\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}$ can be obtained by $\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}=\left\langle\mathrm{Re}(\mathsf{f}_{k})\cos\theta_{i}+\mathrm{Im}(\mathsf{f}_{k})\sin\theta_{i}\right\rangle_{\rho_{i}}$ and $\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}=\left\langle\mathrm{Im}(\mathsf{f}_{k})\cos\theta_{i}-\mathrm{Re}(\mathsf{f}_{k})\sin\theta_{i}\right\rangle_{\rho_{i}}$, respectively. By (15), we can obtain the system of congruences $\mathsf{f}_{k}\equiv\mathsf{y}_{ki}e^{\mathrm{i}\theta_{i}}\mod\mathsf{M}_{i},\ i=1,2,\ldots,L$. For convenience, we let $\mathsf{N}=\mathsf{f}_{k}$, and let $\mathsf{r}_{i}=\mathsf{y}_{ki}e^{\mathrm{i}\theta_{i}}$, then we have $\mathsf{N}\equiv\mathsf{r}_{i}\mod\mathsf{M}_{i},\ i=1,2,\ldots,L$. When there is no error in any remainder, $\mathsf{f}_{k}$ or $\mathsf{N}$ can be recovered by the C-CRT. However, in practical applications, the obtained remainders $\mathsf{r}_{i}$ may have errors. Let the erroneous remainders be $\tilde{\mathsf{r}}_{i}=\mathsf{r}_{i}+\Delta\mathsf{r}_{i}$, where $i=1,2,\ldots,L$, and $\Delta\mathsf{r}_{i}$ represents the error in the $i$-th remainder. In the following, we consider the MLE C-CRT based on the assumption of wrapped complex Gaussian distributions of the errors, which is to estimate the complex value $\mathsf{N}$ from its erroneous complex remainders $\tilde{\mathsf{r}}_{1},\tilde{\mathsf{r}}_{2},\ldots,\tilde{\mathsf{r}}_{L}$.
@@ III-A Circular Distance and Wrapped Distributions
Note that $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=d_{\mathsf{M}}(\mathsf{r},\left\langle\mathsf{N}\right\rangle_{\mathsf{M}})$ by Property 4, we consider $\mathsf{N}\in\mathcal{F}_{\mathsf{M}}$ without loss of generality. Since $\mathsf{r}$, $\mathsf{N}\in\mathcal{F}_{\mathsf{M}}$, we have $\mathsf{r}e^{-\mathrm{i}\theta}$, $\mathsf{N}e^{-\mathrm{i}\theta}\in\mathcal{F}_{\rho}$ by Property 1. Hence, $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right)$, $\mathrm{Im}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right)\in(-\rho,\rho)$. Note that $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=\mathsf{r}-\mathsf{N}-\left[\frac{\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}}{\rho}\right]\mathsf{M}$. That is, $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=\mathsf{r}-\mathsf{N}-\left[\frac{\mathrm{Re}(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta})}{\rho}\right]\mathsf{M}-\mathrm{i}\left[\frac{\mathrm{Im}(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta})}{\rho}\right]\mathsf{M}$. Thus, $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})$ is discontinuous at $\mathsf{r}$ when the real or imaginary part of $\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}$ equals $\frac{\rho}{2}$ or $-\frac{\rho}{2}$. Note that since the proofs of these four cases are similar, we only consider the case when $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right)=\frac{\rho}{2}$, i.e., $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}\right)=\frac{\rho}{2}+\mathrm{Re}\left(\mathsf{N}e^{-\mathrm{i}\theta}\right)$. Since $\mathsf{r}e^{-\mathrm{i}\theta}=(y\sin\theta+x\cos\theta)+(y\cos\theta-x\sin\theta)\mathrm{i}$, we have $y\sin\theta+x\cos\theta=\frac{\rho}{2}+\mathrm{Re}\left(\mathsf{N}e^{-\mathrm{i}\theta}\right)$. Hence, $\mathsf{r}\in\mathcal{D}$. Since $\mathcal{D}$ contains at most four line segments, its measure is zero. ∎
@@
We now consider that the unknown complex number $\mathsf{N}$ to determine is noisy in its observation and the noise follows a complex Gaussian distribution as commonly assumed. This noise results in noises in the remainders. To study the noises in the remainders, we introduce the multivariate wrapped distributions below. Let $f_{\mathbf{X}}(\mathbf{x})$ be the probability density function (pdf) of a 2D random variable $\mathbf{X}$. As described in [40], the pdf of $\mathbf{Y}\equiv\mathbf{X}\mod\mathbf{I}$ is $f_{\mathbf{Y}}(\mathbf{y})=\sum_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{y}+\mathbf{k})$, where $\mathbf{y}\in\mathcal{N}(\mathbf{I})$, and $\mathbf{I}$ is the identity matrix. Next, we consider the pdf of $\mathbf{Z}\equiv\mathbf{X}\mod\mathbf{M}$, where $\mathbf{M}$ is an invertible matrix. Clearly, $\mathbf{M}^{-1}\mathbf{Z}\equiv\mathbf{M}^{-1}\mathbf{X}\mod\mathbf{I}$. If we let $\mathbf{Z}^{\prime}=\mathbf{M}^{-1}\mathbf{Z}$, then $f_{\mathbf{Z}^{\prime}}(\mathbf{z}^{\prime})=|\det(\mathbf{M})|\sum_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{M}(\mathbf{z}^{\prime}+\mathbf{k}))$, where $\det(\mathbf{M})$ is the determinant of $\mathbf{M}$. Consequently, the pdf of $\mathbf{Z}$ is
$f_{\mathbf{Z}}(\mathbf{z})=|\det(\mathbf{M}^{-1})|f_{\mathbf{Z}^{\prime}}(\mathbf{M}^{-1}\mathbf{z})=\sum\limits_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{z}+\mathbf{M}\mathbf{k}),$ (18)
@@
where $\mathbf{z}\in\mathcal{N}(\mathbf{M})$. Then, (18) can be used to represent the pdf of the complex random variable $\mathsf{X}=\mathsf{N}+\mathsf{W}$, where $\mathsf{W}$ is the noise of 00 mean and $\mathsf{N}$ is the true unknown complex number to determine. Specifically, if $\mathsf{W}$ follows a complex Gaussian distribution, then the pdf of $\mathsf{X}$ is $f_{\mathsf{X}}(\mathsf{x})=\frac{1}{2\pi\sigma^{2}}\exp\left\{-\frac{\left|{\mathsf{x}}-\mathsf{N}\right|^{2}}{2\sigma^{2}}\right\}$, where $\sigma^{2}$ is the variance of both the real and imaginary parts of $\mathsf{W}$. Thus, from (18) we have the pdf of $\mathsf{R}\equiv\mathsf{X}\mod\mathsf{M}$:
@@
$\textstyle f_{\mathsf{R}}(\mathsf{r})$ $\textstyle=\frac{1}{2\pi\sigma^{2}}\sum\limits_{\mathsf{k}^{\prime}\in\mathbb{Z}[\mathrm{i}]}\exp\left\{-\frac{|\mathsf{r}-\mathsf{N}-\left[\frac{\mathsf{r}-\mathsf{N}}{\mathsf{M}}\right]\mathsf{M}+\mathsf{k}^{\prime}\mathsf{M}|^{2}}{2\sigma^{2}}\right\}$
$\textstyle=\frac{1}{2\pi\sigma^{2}}\sum\limits_{\mathsf{k}^{\prime}\in\mathbb{Z}[\mathrm{i}]}\exp\left\{-\frac{|d_{\mathsf{M}}(\mathsf{r},\mathsf{N})+\mathsf{k}^{\prime}\mathsf{M}|^{2}}{2\sigma^{2}}\right\}.$ (20)
@@
$\textstyle\iint_{\mathcal{F}_{\mathsf{M}}}\exp\left\{-\frac{|d_{\mathsf{M}}(\mathsf{r},\mathsf{N})|^{2}}{2\sigma^{2}}\right\}\mathrm{d}x\mathrm{d}y$ (22)
$\textstyle=\lim_{n\to\infty}\frac{|\mathsf{M}|^{2}}{n}\sum_{j=1}^{n}\exp\left\{-\frac{u_{j}^{2}+v_{j}^{2}}{2\sigma^{2}}\right\}$
$\textstyle=\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v.$
@@
$\textstyle\Big{(}\int_{-\frac{T}{2}}^{\frac{T}{2}}\exp\Big{\{}-\frac{u^{2}}{2\sigma^{2}}$$\textstyle\Big{\}}\mathrm{d}u\Big{)}^{2}=\iint_{\mathcal{S}_{T}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v$
$\textstyle<\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v.$ (23)
Note that $\frac{1}{\sqrt{2\pi}\sigma}\int_{-3\sigma}^{3\sigma}\exp\left\{-\frac{u^{2}}{2\sigma^{2}}\right\}\mathrm{d}u=0.9973$. Since $|\mathsf{M}|\geq 6\sqrt{2}\sigma$, i.e., $T\geq 6\sigma$, we have
@@
By (III-A), we have $\frac{1}{2\pi\sigma^{2}}\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v>0.9973^{2}$. This leads to (21) by (22) . ∎

### v3

@@ II-B Problem Description and SR-ADCs
Now we consider the problem of recovery of sampled values for complex-valued bandlimited signals using the C-CRT. For a single SR-ADC, one can compute the modulus of real numbers. Two combined SR-ADCs can obtain a modulus sampler with Gaussian integer $\mathsf{M}_{i}$, as illustrated in Fig. 3. For convenience, we denote $\mathsf{M}_{i}=\rho_{i}e^{\mathrm{i}\theta_{i}}$, where $\theta_{i}$ represents the angle of $\mathsf{M}_{i}$, and $\rho_{i}$ represents the dynamic range of the SR-ADC. Let $T$ be the sampling interval length of each SR-ADC and $\mathsf{f}_{k}=\mathsf{f}(kT)$, where $\mathsf{f}(t)$ is a complex-valued bandlimited signal and $k\in\mathbb{Z}$. Then, the output $\mathsf{y}_{ki}$ can be expressed as
@@
By applying a phase shift, one can obtain $\mathsf{y}_{ki}$ through separately applying the modulo operation to its real and imaginary parts. To be specific, if we rewrite $\mathsf{y}_{ki}$ as
$\mathsf{y}_{ki}=\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}+\mathrm{i}\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}},$
then $\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}$ and $\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}$ can be obtained by
$\left\langle\mathrm{Re}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}=\left\langle\mathrm{Re}(\mathsf{f}_{k})\cos\theta_{i}+\mathrm{Im}(\mathsf{f}_{k})\sin\theta_{i}\right\rangle_{\rho_{i}}$
and
$\left\langle\mathrm{Im}(\mathsf{f}_{k}e^{-\mathrm{i}\theta_{i}})\right\rangle_{\rho_{i}}=\left\langle\mathrm{Im}(\mathsf{f}_{k})\cos\theta_{i}-\mathrm{Re}(\mathsf{f}_{k})\sin\theta_{i}\right\rangle_{\rho_{i}},$
respectively. By (16), we can obtain the system of congruences
$\mathsf{f}_{k}\equiv\mathsf{y}_{ki}e^{\mathrm{i}\theta_{i}}\mod\mathsf{M}_{i},\ i=1,2,\ldots,L.$
For convenience, we let $\mathsf{N}=\mathsf{f}_{k}$, and let $\mathsf{r}_{i}=\mathsf{y}_{ki}e^{\mathrm{i}\theta_{i}}$, then we have
$\mathsf{N}\equiv\mathsf{r}_{i}\mod\mathsf{M}_{i},\ i=1,2,\ldots,L.$
When there is no error in any remainder, $\mathsf{f}_{k}$ or $\mathsf{N}$ can be recovered by the C-CRT. However, in practical applications, the obtained remainders $\mathsf{r}_{i}$ may have errors. Let the erroneous remainders be
$\tilde{\mathsf{r}}_{i}=\mathsf{r}_{i}+\Delta\mathsf{r}_{i},$
where $i=1,2,\ldots,L$, and $\Delta\mathsf{r}_{i}$ represents the error in the $i$-th remainder. In the following, we consider the MLE C-CRT based on the assumption of wrapped complex Gaussian distributions of the errors, which is to estimate the complex value $\mathsf{N}$ from its erroneous complex remainders $\tilde{\mathsf{r}}_{1},\tilde{\mathsf{r}}_{2},\ldots,\tilde{\mathsf{r}}_{L}$.
@@ III-A Circular Distance and Wrapped Distributions
Note that $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=d_{\mathsf{M}}(\mathsf{r},\left\langle\mathsf{N}\right\rangle_{\mathsf{M}})$ by Property 4, we consider $\mathsf{N}\in\mathcal{F}_{\mathsf{M}}$ without loss of generality. Since $\mathsf{r},\mathsf{N}\in\mathcal{F}_{\mathsf{M}}$, we have $\mathsf{r}e^{-\mathrm{i}\theta},\mathsf{N}e^{-\mathrm{i}\theta}\in\mathcal{F}_{\rho}$ by Property 1. Hence, $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right),\mathrm{Im}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right)\in(-\rho,\rho)$. Note that
$d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=\mathsf{r}-\mathsf{N}-\left[\frac{\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}}{\rho}\right]\mathsf{M}.$
That is,
$d_{\mathsf{M}}(\mathsf{r},\mathsf{N})=\mathsf{r}-\mathsf{N}-\left[\frac{\mathrm{Re}(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta})}{\rho}\right]\mathsf{M}-\mathrm{i}\left[\frac{\mathrm{Im}(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta})}{\rho}\right]\mathsf{M}.$
Thus, $d_{\mathsf{M}}(\mathsf{r},\mathsf{N})$ is discontinuous at $\mathsf{r}$ when the real or imaginary part of $\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}$ equals $\frac{\rho}{2}$ or $\frac{\rho}{2}$. Note that since the proofs of these four cases are similar, we only consider the case when $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}-\mathsf{N}e^{-\mathrm{i}\theta}\right)=\frac{\rho}{2}$, i.e., $\mathrm{Re}\left(\mathsf{r}e^{-\mathrm{i}\theta}\right)=\frac{\rho}{2}+\mathrm{Re}\left(\mathsf{N}e^{-\mathrm{i}\theta}\right)$. Since
$\mathsf{r}e^{-\mathrm{i}\theta}=(y\sin\theta+x\cos\theta)+(y\cos\theta-x\sin\theta)\mathrm{i},$
we have
$y\sin\theta+x\cos\theta=\frac{\rho}{2}+\mathrm{Re}\left(\mathsf{N}e^{-\mathrm{i}\theta}\right).$
Hence, $\mathsf{r}\in\mathcal{D}$. Since $\mathcal{D}$ contains at most four line segments, its measure is zero. ∎
@@
We now consider that the unknown complex number $\mathsf{N}$ to determine is noisy in its observation and the noise follows a complex Gaussian distribution as commonly assumed. This noise results in noises in the remainders. To study the noises in the remainders, we introduce the multivariate wrapped distributions below. Let $f_{\mathbf{X}}(\mathbf{x})$ be the probability density function (pdf) of a 2D random variable $\mathbf{X}$. As described in [33], the pdf of $\mathbf{Y}\equiv\mathbf{X}\mod\mathbf{I}$ is
$f_{\mathbf{Y}}(\mathbf{y})=\sum_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{y}+\mathbf{k}),$
where $\mathbf{y}\in\mathrm{FPD}(\mathbf{I})$, and $\mathbf{I}$ is the identity matrix. Next, we consider the pdf of $\mathbf{Z}\equiv\mathbf{X}\mod\mathbf{M}$, where $\mathbf{M}$ is an invertible matrix. Clearly,
$\mathbf{M}^{-1}\mathbf{Z}\equiv\mathbf{M}^{-1}\mathbf{X}\mod\mathbf{I}.$
If we let $\mathbf{Z}^{\prime}=\mathbf{M}^{-1}\mathbf{Z}$, then
$f_{\mathbf{Z}^{\prime}}(\mathbf{z}^{\prime})=|\det(\mathbf{M})|\sum_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{M}(\mathbf{z}^{\prime}+\mathbf{k})),$
where $\det(\mathbf{M})$ is the determinant of $\mathbf{M}$. Consequently, the pdf of $\mathbf{Z}$ is
$\hskip 0.0ptf_{\mathbf{Z}}(\mathbf{z})=|\det(\mathbf{M}^{-1})|f_{\mathbf{Z}^{\prime}}(\mathbf{M}^{-1}\mathbf{z})=\sum\limits_{\mathbf{k}\in\mathbb{Z}^{2}}f_{\mathbf{X}}(\mathbf{z}+\mathbf{M}\mathbf{k}),$ (19)
@@
where $\mathbf{z}\in\mathrm{FPD}(\mathbf{M})$. Then, (19) can be used to represent the pdf of the complex random variable $\mathsf{X}=\mathsf{N}+\mathsf{W}$, where $\mathsf{W}$ is the noise of 0 mean and $\mathsf{N}$ is the true unknown complex number to determine. Specifically, if $\mathsf{W}$ follows a complex Gaussian distribution, then the pdf of $\mathsf{X}$ is
$f_{\mathsf{X}}(\mathsf{x})=\frac{1}{2\pi\sigma^{2}}\exp\left\{-\frac{\left|{\mathsf{x}}-\mathsf{N}\right|^{2}}{2\sigma^{2}}\right\},$
where $\sigma^{2}$ is the variance of both the real and imaginary parts of $\mathsf{W}$. Thus, from (19) we have the pdf of $\mathsf{R}\equiv\mathsf{X}\mod\mathsf{M}$:
@@
$f_{\mathsf{R}}(\mathsf{r})$ $=\frac{1}{2\pi\sigma^{2}}\sum\limits_{\mathsf{k}^{\prime}\in\mathbb{Z}[\mathrm{i}]}\exp\left\{-\frac{|\mathsf{r}-\mathsf{N}-\left[\frac{\mathsf{r}-\mathsf{N}}{\mathsf{M}}\right]\mathsf{M}+\mathsf{k}^{\prime}\mathsf{M}|^{2}}{2\sigma^{2}}\right\}$
$=\frac{1}{2\pi\sigma^{2}}\sum\limits_{\mathsf{k}^{\prime}\in\mathbb{Z}[\mathrm{i}]}\exp\left\{-\frac{|d_{\mathsf{M}}(\mathsf{r},\mathsf{N})+\mathsf{k}^{\prime}\mathsf{M}|^{2}}{2\sigma^{2}}\right\}.$ (21)
@@
$\iint_{\mathcal{F}_{\mathsf{M}}}\exp\left\{-\frac{|d_{\mathsf{M}}(\mathsf{r},\mathsf{N})|^{2}}{2\sigma^{2}}\right\}\mathrm{d}x\mathrm{d}y=\lim_{n\to\infty}\frac{|\mathsf{M}|^{2}}{n}\sum_{j=1}^{n}\exp\left\{-\frac{u_{j}^{2}+v_{j}^{2}}{2\sigma^{2}}\right\}=\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v.$ (23)
@@
$\left(\int_{-\frac{T}{2}}^{\frac{T}{2}}\exp\Big{\{}-\frac{u^{2}}{2\sigma^{2}}\Big{\}}\mathrm{d}u\right)^{2}=\iint_{\mathcal{S}_{T}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v<\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v.$ (24)
Note that
$\frac{1}{\sqrt{2\pi}\sigma}\int_{-3\sigma}^{3\sigma}\exp\left\{-\frac{u^{2}}{2\sigma^{2}}\right\}\mathrm{d}u=0.9973.$
Since $|\mathsf{M}|\geq 6\sqrt{2}\sigma$, i.e., $T\geq 6\sigma$, we have
@@
By (24), we have
$\frac{1}{2\pi\sigma^{2}}\iint_{\mathcal{S}_{\mathsf{M}}}\exp\left\{-\frac{u^{2}+v^{2}}{2\sigma^{2}}\right\}\mathrm{d}u\mathrm{d}v>0.9973^{2}.$
This leads to (22) by (23) . ∎